# Evaluating the Quality of the Quantified Uncertainty for (Re)Calibration of Data-Driven Regression Models

### v1

@@ 1 Introduction
An essential component of any recalibration method is a calibration metric to evaluate the accuracy and QQU. Without such a metric, the effect of recalibration cannot be assessed. Yet, here lies a critical issue: numerous metrics exist, but they differ widely in definition, scaling, and interpretation. For instance, the Quantile Calibration Error (QCE) by Küppers et al. (2023) and the Calibration Score (CalS) by Kuleshov et al. (2018) both aim to assess model calibration (recalibration performance) but follow completely different approaches and yield values that are not directly comparable. This may not be problematic when evaluating a single model before and after recalibration, as only the metric difference is of interest. However, it becomes a significant limitation when trying to compare different recalibration methods or benchmark models across datasets, since practitioners may be misled into preferring one method over another based solely on the choice of metric rather than true improvements in calibration.
@@ 3.7 Comparability Between Metrics
When comparing different recalibration approaches, the preferred method is usually the one that produces the largest improvement in the chosen metric. This practice assumes that all metrics respond consistently to improvements in model calibration. However, a closer look at the metrics reveals significant differences in their formulation. For example, the CWC depends on a user-defined hyperparameter, the CRPS has the same unit as the target variable, and the NLL is influenced by the logarithm of the predicted variance. Thus, beyond differences in scale, these metrics also vary in how strongly they weight accuracy versus QQU.
@@ 4 Benchmark
To evaluate the consistency of different metrics for assessing model calibration, we conduct a comprehensive benchmark. A diverse selection of real-world and synthetic datasets is used to train data-driven models for predicting the respective target variable. For each of the models the performance is then assessed using a variety of calibration metrics. The results reveal significant inconsistencies between metrics: for the same model, one metric may indicate high-quality uncertainty estimates, while another deems the same model’s uncertainty poor.
@@
To evaluate the consistency of the metrics in the first two parts of the benchmark, we introduce an evaluation scheme based on the comparison of the normalized model performance. As outlined previously, the metrics differ significantly in scale and interpretation – some being bounded, others semi-bounded (e.g., from 0 to $\infty$) or unbounded. This makes direct comparison of raw metric values across datasets unreliable. Furthermore, many metrics originate from recalibration literature, where relative changes (e.g., before vs. after recalibration) are the primary focus. To address these issues, we compute the metric values and normalize them by division with the mean value of said metric across all datasets. This preserves relative performance differences while bringing all metrics onto a comparable scale. For the PCIP we consider the absolute difference to the nominal confidence level. We then sort the datasets by performance and visualize the resulting values as line plots. This is inspired by the fact that a good model should be deemed good by all metrics. Ideally, consistent metrics should exhibit similar trends – that is, their lines should follow the same general order across datasets. Strong deviations in direction or ordering (for instance, if a model performs well under one metric but poorly under another) indicate inconsistency and cast doubt on the reliability of the metric for evaluation of a model.
Motivated by these findings, the third part of the benchmark follows a more classical setup: we compare each model before and after recalibration using all metrics. The goal is not to determine whether recalibration improves model performance in general, but rather to illustrate how metric inconsistency affects the interpretation of recalibration results. As the results demonstrate, the perceived effectiveness of recalibration depends heavily on the chosen metric – highlighting how easily recalibration outcomes can be framed positively or negatively by selectively reporting favorable metrics.
@@
The results confirm that substantial inconsistencies exist between the metrics. In the case of synthetic datasets and artificial miscalibration these inconsistencies persist, suggesting that the disagreement stems from the metrics themselves rather than from data-related noise or anomalies. However, we show that the CWC and ENCE come out superior in our benchmark.
@@ 4.1 Case study: real-world data
The results of the real-world dataset evaluation are presented in Fig. 2. The behavior of the metrics reveals two distinct groups that show internal agreement in how they assess model performance across the datasets.
@@
The first and largest group – located in the upper left figure – consists of RMSE, Sharpness, NLL, CRPS, IS, and Pinball Loss. These metrics produce similar orderings of model performance and follow consistent trends across datasets. A commonality among them is that they are proper scoring rules and do not rely on discretizing the prediction or target space into bins. Instead, they directly assess the predictive distribution as a whole, including accuracy and uncertainty, leading to smoother and more globally consistent behavior. This theoretical foundation likely contributes to their empirical agreement.
@@
A second distinct cluster comprises CalS, ECPE and QCE. These metrics show strong agreement in their ordering of model performance, which can likely be attributed to their shared structural basis: all three rely on threshold-based and binned evaluation. CalS and ECPE assess model quality across multiple confidence levels, measuring whether predicted intervals correctly cover the target variable. QCE, while based on quantiles rather than confidence intervals, follows a similar logic by comparing predicted versus empirical quantiles across a spectrum of thresholds. This localized, thresholded evaluation approach distinguishes them from the first group of metrics and likely contributes to their internal consistency.
@@
The remaining metrics – Prediction Interval Coverage Probability (PICP), Coverage Width Criterion (CWC), Expected Normalized Calibration Error (ENCE), and Uncertainty Calibration Error (UCE) – exhibit less consistent behavior. While there are partial overlaps or isolated agreements with the two main groups, they do not form a cohesive cluster.
@@
Importantly, even within the internally consistent groups, disagreements between groups are substantial. For instance, in the case of the first dataset, the proper scoring rules (Group 1) consistently rate the ensemble model as highly reliable, while the threshold-based calibration metrics (Group 2) assign it the lowest performance. Such a reversal in evaluation highlights how the choice of metric can drastically influence conclusions, even when applied to the same model and data.
@@
Figure 2: Normalized metric values for the deep ensembles across 16 real-world datasets. The values are normalized per metric by their mean across datasets. Group 1 consists of proper scoring rules and closely related metrics that aggregate over the full predictive distribution without discretizing it. Group 2 shows internal agreement likely due to their shared use of threshold-based evaluations across confidence levels or quantiles. Group 3 shows less/no internal consistency. Notably, metrics from different groups frequently disagree on which model performs best on a given dataset, underscoring the substantial inconsistency in uncertainty evaluation. The IDs to the respective datasets are provided in Appendix A.
@@ 4.2 Case study: synthetic data
The results on the synthetic datasets (see Fig. 3) show a clustering of metrics identical to that observed for the real-world data. Again, proper scoring rules and related metrics (Group 1) behave consistently among themselves, as do the threshold-based calibration metrics (Group 2). This confirms that the inconsistencies between metric groups are not caused by data noise or aleatoric uncertainty, since the synthetic datasets are specifically constructed to be noise-free and well-behaved. Instead, the observed disagreements appear to stem from fundamental differences in what the metrics evaluate, reinforcing the findings from the real-world benchmark.
Figure 3: Normalized metric values for the deep ensembles across 10 synthetic datasets. The values are normalized per metric by their mean across datasets. The same three metric groups emerge as in the real-world setting: Group 1 includes proper scoring rules or closely related metrics that aggregate over the full predictive distribution without discretization. Group 2 consists of metrics with threshold-based evaluations across confidence levels or quantiles. Group 3 shows limited internal agreement. The persistence of these groups, despite the absence of aleatoric noise in the synthetic datasets, suggests that the inconsistencies are due to intrinsic differences in metric definitions. This emphasizes the need for caution when selecting metrics to evaluate uncertainty quality. The IDs to the respective datasets are provided in Appendix A.
The third part of the benchmark further demonstrates that these inconsistencies have a significant impact on the evaluation of recalibration methods, as the perceived effectiveness of recalibration varies strongly depending on the chosen metric.
@@ 4.3 Case study: recalibration
Crucially, for nearly every dataset and recalibration method, at least one metric reports an improvement while another indicates a deterioration. This inconsistency highlights a serious challenge: the effectiveness of a recalibration method can appear entirely different depending on which calibration metric is used. Furthermore, a non-negligible number of recalibrations even lead to a decrease in model performance across several metrics, suggesting that recalibration is not universally beneficial in practice – although evaluating this properly is complicated by the very inconsistencies demonstrated here and the lack of ground truth uncertainties.
@@ 4.4 Case study: controlled miscalibration
As in the recalibration case study (Section 4.3), we focus on changes in metric values exceeding a relative threshold of 3%. Fig. 5 shows how each metric responds to the different miscalibration types. Metrics in Group 2 (threshold-based metrics) consistently detect the constant standard deviation offset, while metrics in Group 1 (proper scoring rules) generally do not. The heterogeneous standard deviation offset yields less consistent results, with some Group 2 metrics misidentifying the degraded data as better calibrated, highlighting their sensitivity but also potential unreliability. The constant mean prediction offset scenario is particularly revealing: among the proper scoring rules, only the NLL reliably flags the miscalibration, while others misinterpret the degraded predictions as improved. In contrast, most metrics from Groups 2 and 3 correctly identify the calibration deterioration.
The fourth and most complex scenario with simultaneous heterogeneous offset of both mean and standard deviation exposes further nuances. While NLL again stands out within Group 1 for its consistent detection, other proper scoring rules show limited sensitivity. Group 2 metrics detect changes in calibration across the board, but about half erroneously indicate an improvement, suggesting poor alignment with ground-truth reliability. Notably, metrics such as CWC, ENCE, and UCE which previously did not cluster clearly with other groups emerge as the most consistent, correctly identifying degradation in nearly all datasets and scenarios.
@@
Figure 5: Relative change in metric values after artificially miscalibrating predictions across four scenarios. Each heatmap cell indicates whether a metric detects an improvement (blue), degradation (orange), or negligible change (grey). Metrics are colored by their previously identified group: purple (Group 1), teal (Group 2), black (Group 3).
@@
Because some real-world datasets are small (fewer than 200 samples), the random sampling of predictions and binning used in some metrics can introduce high variability. This results in instability across benchmark runs. To account for this, we repeat the entire miscalibration benchmark 100 times and count how often each metric correctly detects degradation (defined as a $\geq 3\%$ worse calibration score) in the fourth, most realistic scenario.
@@
Fig. 6 summarizes these results. The ability to detect miscalibration varies significantly between datasets. Datasets 3 and 9, which have the fewest samples, exhibit the lowest detection frequencies – highlighting the impact of sample size on calibration evaluation reliability. As dataset size increases, so does the likelihood of correct detection. Among the proper scoring rules, only NLL reliably flags the miscalibration. Group 2 metrics behave similarly (group internal) but remain sensitive to noise in smaller datasets.
CWC, ENCE, and UCE again stand out for their robustness. Despite lacking clear internal consistency across previous benchmark parts, they consistently detect miscalibration in this setting and outperform Group 2. The ENCE is marginally more reliable than UCE, while CWC appears to be the most stable across datasets and runs.
@@ 4.5 Recommendations
Our results suggest that calibration metrics can reliably detect miscalibration effects when evaluated on datasets with at least 500–1000 samples, assuming typical number of bins for binning-based approaches (e.g., 10 bins). For smaller datasets, stochastic effects (such as sampling variability and binning noise) can obscure true calibration quality.
Among all evaluated metrics, three stand out as particularly informative.
@@
The ENCE also performed well, demonstrating consistent sensitivity to miscalibration in both mean and uncertainty, while not requiring any hyperparameters. Unlike the CWC, ENCE evaluates calibration across multiple bins and is thus better suited to detect local patterns of miscalibration. It also avoids the ambiguity of choosing a specific confidence level and can provide a more fine-grained view of calibration quality.
@@ 5 Discussion
We have introduced a taxonomy that describes calibration metrics as metrics that capture the QQU and accuracy of a model. In theory, for most recalibration procedures, capturing the QQU alone would suffice, but potential influences on accuracy would then not be captured. We therefore consider the combination of the two performance indicators in a calibration metric to be expedient. This is particularly relevant in view of potential future recalibration procedures that influence both.
@@
Another important consideration is that we did not assess the initial calibration state of models prior to recalibration and thus might have tried to recalibrate an already calibrated model. As the metric values are often non-interpretable on their own, this is a common circumstance. For the purpose of comparing metric behaviors, this is inconsequential: regardless of whether recalibration improves the model, the inconsistency in metric response remains.
@@
Finally, the set of metrics considered in this study is not exhaustive. For instance, Kristoffersson Lind et al. (2024) explored metrics such as Spearman rank correlation, but deemed it inferior to CalS, which is why we refrained from including it.

### v2

@@ 1 Introduction
An essential component of any recalibration method is a calibration metric to evaluate the accuracy and QQU. Without such a metric, the effect of recalibration cannot be assessed. Yet, here lies a critical issue: numerous metrics exist, but they differ widely in definition, scaling, and interpretation. For instance, the Quantile Calibration Error (QCE) by Küppers et al. (2023) and the Calibration Score (CalS) by Kuleshov et al. (2018) both aim to assess model calibration (recalibration performance) but follow completely different approaches and yield values that are not directly comparable. This may not be problematic when evaluating a single model before and after recalibration, as only the metric difference is of interest. However, it becomes a considerable limitation when trying to compare different recalibration methods or benchmark models across datasets, since practitioners may be misled into preferring one method over another based solely on the choice of metric rather than true improvements in calibration.
@@ 3.7 Comparability Between Metrics
When comparing different recalibration approaches, the preferred method is usually the one that produces the largest improvement in the chosen metric. This practice assumes that all metrics respond consistently to improvements in model calibration. However, a closer look at the metrics reveals substantial differences in their formulation. For example, the CWC depends on a user-defined hyperparameter, the CRPS has the same unit as the target variable, and the NLL is influenced by the logarithm of the predicted variance. Thus, beyond differences in scale, these metrics also vary in how strongly they weight accuracy versus QQU.
@@ 4 Benchmark
To evaluate the consistency of different metrics for assessing model calibration, we conduct a comprehensive benchmark. A diverse selection of real-world and synthetic datasets is used to train data-driven models for predicting the respective target variable. For each of the models the performance is then assessed using a variety of calibration metrics. The results reveal substantial inconsistencies between metrics: for the same model, one metric may indicate high-quality uncertainty estimates, while another deems the same model’s uncertainty poor.
@@
To evaluate the consistency of the metrics in the first two parts of the benchmark, we introduce an evaluation scheme based on the comparison of the normalized model performance. As outlined previously, the metrics differ substantially in scale and interpretation – some being bounded, others semi-bounded (e.g., from 0 to $\infty$) or unbounded. This makes direct comparison of raw metric values across datasets unreliable. Furthermore, many metrics originate from recalibration literature, where relative changes (e.g., before vs. after recalibration) are the primary focus. To address these issues, we compute the metric values and normalize them by division with the mean value of said metric across all datasets. This preserves relative performance differences while bringing all metrics onto a comparable scale. For the PCIP we consider the absolute difference to the nominal confidence level. We then sort the datasets by performance and visualize the resulting values as line plots. This is inspired by the fact that a good model should be deemed good by all metrics. Ideally, consistent metrics should exhibit similar trends – that is, their lines should follow the same general order across datasets. Strong deviations in direction or ordering (for instance, if a model performs well under one metric but poorly under another) indicate inconsistency and cast doubt on the reliability of the metric for evaluation of a model. In addition, we rank the datasets for each metric and assess the Spearman’s rank correlation coefficient as consistent metrics should yield a high rank correlation. Both methods attest severe inconsistencies between metrics.
Motivated by these findings, the third part of the benchmark follows a more classical setup: we compare each model before and after recalibration using all metrics. The goal is not to determine whether recalibration improves model performance in general, but rather to illustrate how metric inconsistency affects the interpretation of recalibration results. The results demonstrate that the perceived effectiveness of recalibration depends heavily on the chosen metric – highlighting how easily recalibration outcomes can be framed positively or negatively by selectively reporting favorable metrics.
@@
Overall, results confirm that substantial inconsistencies exist between the metrics. In the case of synthetic datasets and artificial miscalibration these inconsistencies persist, suggesting that the disagreement stems from the metrics themselves rather than from data-related noise or anomalies. However, we show that the CWC and ENCE come out superior in our benchmark.
@@ 4.1 Case study: real-world data
The Spearman’s rank correlation coefficients between the metrics’ ranking of the datasets are presented as a heatmap in Fig. 2. Comparing the correlation reveals two distinct groups that show internal agreement in how they assess model performance across the datasets.
Figure 2: Heatmap of the Spearman’s rank correlation coefficients using the real-world data. The evaluation metrics label color indicates the identified metric clusters: Purple for proper scoring rules and related metrics (Group 1), teal for threshold-based metrics (Group 2), and black for others (Group 3).
@@
The first and largest group – located in the upper left corner – consists of RMSE, Sharpness, NLL, CRPS, IS, and Pinball Loss. These metrics produce similar orderings of model performance and follow consistent trends across datasets. A commonality among them is that they are proper scoring rules and do not rely on discretizing the prediction or target space into bins. Instead, they directly assess the predictive distribution as a whole, including accuracy and uncertainty, leading to smoother and more globally consistent behavior. This theoretical foundation likely contributes to their empirical agreement.
@@
A second distinct cluster comprises CalS, ECPE, QCE and ENCE. These metrics show strong agreement in their ordering of model performance, which can likely be attributed to their shared structural basis: all four rely on threshold-based and binned evaluation. CalS and ECPE assess model quality across multiple confidence levels, measuring whether predicted intervals correctly cover the target variable. QCE, while based on quantiles rather than confidence intervals, follows a similar logic by comparing predicted versus empirical quantiles across a spectrum of thresholds. The ENCE discretizes the target domain based on the standard deviations of the predictions of the model. This localized, thresholded evaluation approach distinguishes them from the first group of metrics and likely contributes to their internal consistency. However, it can be seen that the ENCE does not fully agree with the other metrics and the correlation is weaker.
@@
The remaining metrics – PICP, CWC, and UCE – exhibit less consistent behavior. While there are partial overlaps or isolated agreements with the first group, they do not form a cohesive cluster.
Importantly, even within the internally consistent groups, disagreements between groups are substantial. The contrasting ranking pattern between groups – ranging from strong positive correlations to moderate negative correlations – indicates that the order of datasets (from good to bad performance) is partially reversed.
@@
To gain more insights, we visually display the normalized metric values across the 16 real-world datasets in the identified groups (see Fig. 3). Overall, the visual impression supports the correlation and the contrasting ranking pattern between groups can be identified. For instance, in the case of the first dataset, the proper scoring rules (Group 1) consistently rate the ensemble model as highly reliable, while the threshold-based calibration metrics (Group 2) assign it the lowest performance. Such a reversal in evaluation highlights how the choice of metric can drastically influence conclusions, even when applied to the same model and data.
Figure 3: Normalized metric values for the deep ensembles across 16 real-world datasets. The values are normalized per metric by their mean across datasets. Groups 1 and 2 are formed based on similarity due to high correlation. Group 3 shows less/no internal consistency. Notably, metrics from different groups disagree on which model performs best on a given dataset, underscoring the substantial inconsistency in uncertainty evaluation. The IDs to the respective datasets are provided in Appendix A.
@@ 4.2 Case study: synthetic data
The Spearman’s rank correlation heatmap on the synthetic datasets (see Fig. 4) show a similar clustering of metrics to that observed for the real-world data, albeit the contrasting ranking between groups has partly diminished. Most metrics now exhibit strong to no correlation, but no negative values – apart from the ENCE. Nevertheless, the groups still exist. The ENCE, however, shows weak to moderate negative correlations with most other metrics, thereby ranking model performance in the opposite order.
Figure 4: Heatmap of the Spearman’s rank correlation coefficients using synthetic data. The evaluation metrics label color indicates the previously identified metric cluster: Purple for proper scoring rules and related metrics (Group 1), teal for threshold-based metrics (Group 2), and black for others (Group 3).
The visual inspection of the metric values (see Fig. 5) confirms the clustering of metrics – again with the exception of the ENCE.
Figure 5: Normalized metric values for the deep ensembles across 10 synthetic datasets. The values are normalized per metric by their mean across datasets. The persistence of the three groups, despite the absence of aleatoric noise in the synthetic datasets, suggests that the inconsistencies are due to intrinsic differences in metric definitions. This emphasizes the need for caution when selecting metrics to evaluate uncertainty quality. The IDs to the respective datasets are provided in Appendix A.
Overall, the benchmark indicates that the inconsistencies between metric groups are not caused by data noise or aleatoric uncertainty, since the synthetic datasets are specifically constructed to be noise-free and well-behaved. Instead, the observed disagreements appear to stem from fundamental differences in what the metrics evaluate, reinforcing the findings from the real-world benchmark.
The third part of the benchmark further demonstrates that these inconsistencies have a major impact on the evaluation of recalibration methods, as the perceived effectiveness of recalibration varies strongly depending on the chosen metric.
@@ 4.3 Case study: recalibration
Crucially, for nearly every dataset and recalibration method, at least one metric reports an improvement while another indicates a deterioration. This inconsistency highlights a serious challenge: the effectiveness of a recalibration method can appear entirely different depending on which calibration metric is used. Furthermore, a non-negligible number of recalibrations even lead to a decrease in model performance across several metrics, suggesting that recalibration is not universally beneficial in practice – although evaluating this properly is complicated by the very inconsistencies demonstrated here and the lack of ground truth uncertainties. In the following, we therefore conduct a test with controlled miscalibration, where the ground truth is known.
@@ 4.4 Case study: controlled miscalibration
As in the recalibration case study (Section 4.3), we focus on changes in metric values exceeding a relative threshold of 3%. Fig. 7 shows how each metric responds to the different miscalibration types. Metrics in Group 2 (threshold-based metrics) almost consistently detect the constant $\hat{\sigma}$ offset, while metrics in Group 1 (proper scoring rules) generally do not. The heterogeneous $\hat{\sigma}$ offset yields less consistent results, with some Group 2 metrics misidentifying the degraded data as better calibrated, highlighting their sensitivity but also potential unreliability. Here we can see once again the outstanding performance of ENCE in group 2, which usually correctly detects degradation. The constant $\hat{y}$ offset scenario is particularly revealing: among the proper scoring rules, only the NLL reliably flags the miscalibration, while others misinterpret the degraded predictions as improved. In contrast, most metrics from Groups 2 and 3 correctly identify the calibration deterioration.
@@
Figure 7: Relative change in metric values after artificially miscalibrating predictions across four scenarios. Each heatmap cell indicates whether a metric detects an improvement (blue), degradation (orange), or negligible change (grey). It would be correct if a degradation is identified. Metrics are colored by their previously identified group: purple (Group 1), teal (Group 2), black (Group 3).
The fourth and most complex scenario with simultaneous heterogeneous offset of both mean and standard deviation exposes further nuances. While NLL again stands out within Group 1 for its consistent detection, other proper scoring rules show limited sensitivity. Group 2 metrics detect changes in calibration across the board, but for about half of the datasets erroneously an improvement is detected, suggesting poor alignment with ground-truth reliability. Notably, the ENCE quite consistently detects the miscalibration together with metrics such as CWC and UCE – which previously did not cluster clearly with other groups – and thus emerge as the most consistent, correctly identifying degradation in nearly all datasets and scenarios.
@@
Because some real-world datasets are small (fewer than 200 samples), the random sampling of predictions and binning used in some metrics can introduce high variability. This results in instability across benchmark runs. To account for this, we repeat the entire miscalibration benchmark 100 times and count how often each metric correctly detects degradation (defined as a $\geq 3\%$ worse calibration score) in the fourth, most realistic, scenario.
@@
Fig. 8 summarizes these results. The ability to detect miscalibration varies substantially between datasets. For example, datasets 3 and 9, which have the fewest samples, exhibit the lowest detection frequencies – highlighting the impact of sample size on calibration evaluation reliability. As dataset size increases, so does the likelihood of correct detection. Among the proper scoring rules, only NLL reliably flags the miscalibration. Group 2 metrics behave similarly (group internal) but remain sensitive to noise in smaller datasets.
@@
Overall, the CWC, ENCE, and UCE again stand out for their robustness. Despite lacking clear internal consistency across previous benchmark parts, they consistently detect miscalibration in this setting and outperform Group 2. The ENCE is marginally more reliable than UCE, while CWC appears to be the most stable across datasets and runs.
@@ 4.5 Recommendations
Our results suggest that calibration metrics provide consistent results when evaluated on datasets with at least 500–1000 samples, assuming a typical number of bins for binning-based approaches (e.g., 10 bins). For smaller datasets, stochastic effects (such as sampling variability and binning noise) can lead to unstable or contradictory results, regardless of whether these conclusions reflect the true calibration quality.
Among all evaluated metrics, the NLL, CWC and ENCE stand out as particularly informative.
@@
The ENCE performed well, demonstrating consistent sensitivity to miscalibration in both mean and uncertainty, while not requiring any hyperparameters. Unlike the CWC, ENCE evaluates calibration across multiple bins and is thus better suited to detect local patterns of miscalibration. It also avoids the ambiguity of choosing a specific confidence level and can provide a more fine-grained view of calibration quality. However, for small datasets, the ENCE does not always correctly detect miscalibration. Nevertheless, it still outperforms the other metrics of the benchmark.
@@ 5 Discussion
We have introduced a taxonomy that describes calibration metrics as metrics that capture the QQU and accuracy of a model. In theory, for most recalibration procedures, capturing the QQU alone would suffice, but potential influences on accuracy would then not be captured. We therefore consider the combination of the two performance indicators in a calibration metric to be appropriate. This is particularly relevant in view of potential future recalibration procedures that influence both.
@@
Another important consideration is that we did not assess the initial calibration state of models prior to recalibration and thus might have tried to recalibrate an already calibrated model. As the metric values are often non-interpretable on their own, this is a common circumstance. For the purpose of comparing metric behaviors, this is inconsequential: regardless of whether recalibration improves the predictive quality of the model, the inconsistency in the responses of the metrics remain.
@@
Finally, the set of metrics considered in this study is not exhaustive. For instance, Kristoffersson Lind et al. (2024) explored metrics such as Spearman rank correlation between the predicted uncertainty and the error, but deemed it inferior to CalS, which is why we refrained from including it.